INAF queue observing, 2018 September 25-26

Observer: R. Carini, A. Gargiulo
Telescope Operator: D. G. Huerta
Support Astronomer: B. Rothberg
Instrument: LUCI

Summary

At the sunset the sky is mostly clear
instruments: MODS2 red grating is not working
LBC-blue has some issues, LUCI1 works only if the guide star is on the right of the patrol field, LUCI1 AO doesn't work.

Night duration (12\xB0 el twilight) 2:30- 12:15 UT = 9 h 45m

Observing time: 1h 27m (0%)

Weather time loss: 5h 17m (100%)

technical time loss 3h 00m (%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
dard Nisini        
arc flat Nisini      
telluric Perna   LUCI1+LUCI2      

Details

Some issues with the connection between the telescope and LUCI
  • 02:46 the problem has been fixed, the TO is working on the checkout of the AO system.
  • 03:37 we could stars but clouds are coming, we have to close!!!

calibration

Nisini field3
  • dark luci1.20180926.0010-0014, luci2.20180926.0006-0010=
  • flats H luci1.20180926.0015-0019, luci2.20180926.0011-0015=
  • arc H luci1.20180926.0020-0021, luci2.20180926.0016-0017= Ar
  • arc H luci1.20180926.0022-0023, luci2.20180926.0018-0019= Xe
  • dark luci1.0024-0028, luci2.0020-0024=
  • flats J luci1.0029-0033, luci2.0025-0029=
  • arc J luci1.0034-0035, luci2.0030-0031= Ar
  • arc J luci1.0036-0037, luci2.0032-0033= Xe
  • dark luci1.0038-0042, luci2.0034-0038=
  • flats K luci1.0043-0047, luci2.0039-0043=
  • arc K luci1.0048-0049, luci2.0044-0045= Ne
  • arc K luci1.0050-0051, luci2.0046-0047= Ar

08:54 the sky is clearer, TO opens the dome

09:10 start with Nardini program, the telescope can not find the guide star, there is a strange pattern in the detector.

10:00 the problem resolved, TO start the collimation

10:15 the collimation finished,start the telluric HIP1317
  • 10:16 error during commit
  • 10:18 send again the script * error again. the target is to close to the moon.

10:20 slew to Nisini program (seeing on DIMM 1.2''), telluric Hip 17971
  • error in the insert of mask in LUCI2
  • 10:30 mask problem fixed, but new problems with the pointing are happening

10:52 send again the script
  • 10:53 error again

11:01 send again the script
  • the background is too bright, guide star is too faint (14.75 mag in R)/near the moon. We can chose only guide stars on the left of the patrol field and there are not stars brigther than this one.
*11:17 we slew to Perna program, even if the seeing is not perfect
  • load the script for the telluric
  • error, the left side don't see the the star
  • 11:30 : send again
  • error the preset
  • 11:40 send again the script
  • 11:45 filter error (LUCI2)
  • 11:46 we continue the acquisition
  • acq luci1.0053-0055, luci2.0049-0051
  • sci luci1.0056-0059,luci2.0052-0055
  • seeing 1.0''

11:55 send the science script, the guide star is on the left of the patrol field, we change it
  • 12:01 send again the script, but LUCi1 failed, the star for this condition is to faint
  • 12:07 send again with another star but the preset failed (R=14.5)
  • on the right side of the patrol field there are not stars brighter than this one, we can not rotate the field because we are in MOS mode.

-- %USERSIG{RobertaCarini - 2018-09-26}%

Comments

 
Topic revision: r7 - 01 Oct 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback